# Robust Signal Detection with Quadratically Convex Orthosymmetric Constraints

### v2

@@ 11  Set $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$
Set $\left\lVert\nu\right\rVert_{2}=0$ in Corollary A.14, we have
$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)(\widetilde{Y}_{i}-\nu)^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq C_{1}\left(\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right).$ (B.18)
By normalization and standard concentration tools (see A.6), we have
$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)^{\top}\nu\right\rVert_{2}=\left\lVert\nu\right\rVert_{2}\cdot\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)\right\rVert_{2}=C_{0}\cdot\sqrt{Nk^{\star}}\cdot\left\lVert\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\widetilde{Y}_{i}-\nu}{\sqrt{k^{\star}}}\right\rVert_{2}\leq c_{2}\sqrt{Nk^{\star}}$ (B.19)
with (high) constant probability. Now combine (B.18) and (B.19) to get
@@
Bound on II. Since upon the finish the algorithm, we have $\left\lVert\mathbf{1}-\omega\right\rVert_{1}\leq 2\epsilon N$, by Corollary A.16, we know
@@
Bound on III. Since $\left|S\cup C\right|\leq 2\epsilon N$, by Corollary A.16 and an argument similar to the one in II, we also have
@@
$\left\lVert\sum\limits_{i\in[N]\backslash(S\cup C)}\omega_{i}Y_{i}^{\top}Y_{i}-\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq 4c\gamma_{2}.$ (B.20)
Now combine (B.17) and (B.20) and use the triangular inequality, we find that
@@
Again, by choosing the constant in the definition of $\gamma_{2}$ (B.13) sufficiently large, we can make $c$ above sufficiently small. Therefore, we can assume $c<\frac{1}{4}$, which leads to $\left\lVert\sum\limits_{i=1}^{N}\omega_{i}Y_{i}^{\top}Y_{i}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}>\gamma_{2}$, a contradiction of the control condition of Algorithm 4. The proof is completed. ∎
@@
$\gamma_{3}=c\left(\sqrt{Nk^{\star}}+\sqrt{k^{\star}\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right)$ (B.21)
@@
$\omega_{i}=\left(1-\frac{\tau_{i}}{\max\limits_{i}\tau_{i}}\right)\omega_{i},1\leq i\leq N.$ (B.22)
@@
1Set $\gamma_{3}$ as (B.21), and $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
@@ 4  Compute $\tau_{i}=\frac{v_{i}^{2}}{\omega_{i}}\mathbf{1}_{\left\{w_{i}>0\right\}}$
Proof of (i) is same as the one in the proof of Lemma B.12.
@@
Now assume $\omega^{(t)}\in\Lambda_{2}$ and we are still in the iteration, we will prove $\omega^{(t+1)}\in\Lambda_{2}$. By the definition, we only need to prove $5\sum\limits_{i\in C}v_{i}^{2}\geq\sum\limits_{i\in[N]\backslash C}v_{i}^{2}$. Suppose this does not hold, then $\left\lVert v_{C}\right\rVert_{2}^{2}<\frac{1}{6}$. By Corollary A.14 (set $\left\lVert\nu\right\rVert_{2}=0$), we have $\left\lVert\sqrt{D(\omega_{[N]\backslash C}^{(t)})}Y_{[N]\backslash C}Y_{[N]\backslash C}^{{}^{\top}}\sqrt{D(\omega_{[N]\backslash C}^{(t)})}-D(\omega_{[N]\backslash C}^{(t)})\right\rVert_{2}\leq c\gamma_{3}\leq c\lambda^{(t)}$ with high probability. However, since $v$ is the eigenvector corresponding to the maximal eigenvalue, we have
@@
Noting that $\left\lVert v_{C}\right\rVert_{2}^{2}+\left\lVert v_{[N]\backslash C}\right\rVert_{2}^{2}=1$, define $f_{c}(x)=c+(1-c)x+2\sqrt{x(1-x)},0\leq x<\frac{1}{6}$. Note that by selecting the constant in (B.21) sufficiently large, we can let $c\rightarrow 0$. It can be proved that in this case $0\leq f_{0}(x)<1$, which is a contradiction. (ii) is proved.
@@
After Algorithms 4 or 5, we have filtered the samples to ensure the second condition of $\omega$-regularity. Now we turn to the check of the final condition.
Actually, the weight $\omega$ output by Algorithms 4 or 5 already provides us with some useful insight.
Lemma B.15.
Set $\beta_{1}$, $\beta_{2}$ according to (B.11) and $\gamma_{i},i\in\left\{2,3\right\}$ according to (B.13) or (B.21) depending on whether $N\geq k^{\star}$ or not. For the weight $\omega$ output by the Algorithms 4 or 5 and any S[$\CS\subset[N]\backslash$ C with $\left|S\right|\leq\epsilon N$, we have
@@
By Lemma B.8, with high probability the original samples $\widetilde{\mathbf{Y}}$ are $(2\epsilon,\beta_{1},\beta_{2})$-regular. By spliting $N]=([N]\backslash C)\cup C$, we have
@@
Bound on $\text{I}_{1}$. Because $\left|S\right|\leq\epsilon N$, from $\omega$-regularity and Corollary B.10, we know $\text{I}_{1}\leq c_{1}\sqrt{N}\beta_{1}$.
Bound on $\text{I}_{2}$. Because $\omega$ is output by the Algorithms 4 or 5, we have $\left\lVert 1-\sqrt{\omega}\right\rVert_{1}\leq\left\lVert 1-\omega\right\rVert_{1}\leq\epsilon N$. By Corollary B.10 again, we have $\text{I}_{2}\leq c_{2}\beta_{2}$.
Bound on II. From Corollary B.10 and the fact $S\cap C=\varnothing$, we know $\text{II}\leq c_{3}\beta_{2}$.
Bound on III. From Lemma B.12, B.14, and the polarization equality, we have $\text{III}\leq c_{4}\epsilon N\gamma$.
@@
We are now prepared for the final algorithm, which is designed to ensure the third condition of $\omega$-regularity. Given samples $S$ and weight $\omega$ by the algorithms 4 or 5, we compute $\tau_{i}=\left|\left\langle\sqrt{\omega_{i}}Y_{i},\sum\limits_{j=1}^{N}\sqrt{\omega_{j}}Y_{j}\right\rangle-\omega_{i}k^{\star}\right|\mathbf{1}_{\left\{\omega_{i}>0\right\}}$. Then, we set $\omega_{i}$ to zero for the indices $i$ corresponding to the $\epsilon N$ maximal $\tau_{i}$.
@@
This lemma is very similar to Lemma B.15 but stronger in the sense that we allow any $S\subset[N]$ rather than S[$\CS\subset[N]\backslash$ C. This advantage is exactly granted by the sorting and pruning processes in Algorithm 6.
@@
The proof of this lemma is not hard but depends on the following fact. Because we have proved in Lemma B.15 that for any S[$\CS\subset[N]\backslash$ C and $\left|S\right|\leq\epsilon N$, we have $\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma)$, we know that this bound also holds for any $S\subset[N]\backslash\left\{1,2,\dots,\epsilon N\right\}$. With the proning on $\omega$ to remove the first $\epsilon N$ coordinates, we can assume without loss of generality that $S\subset[N]\backslash\left\{1,2,\dots\epsilon N\right\}$. We have
@@
We have finished the proofs of Lemma B.11, B.12, B.14, and B.16, which means we have verified the correctness of $\omega$-regularity for our samples with high probability. The final step is to exploit the property of $\omega$-regularity to distinguish between $H_{0}$ and $H_{1}$.
Lemma B.17.
Let $\omega$ be the weights after running through the algorithm 3, 4 or 5, and 6, and $\beta_{1},\beta_{2}$ seleced as in Lemma B.8. Then, with high probability we have
$\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c\left(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma+\underbrace{\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}+\sqrt{Nk^{\star}}\right)\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)}_{:=\text{R}}\right).$ (B.23)
@@
Bound on I. Define the vector $r\in\mathbb{R}^{N}$ where $r_{i}=1$ for $i\in C$ and $1-\sqrt{\omega_{i}}$ for i[$\Ci\in[N]\backslash$ C. By Algorithm 4 or 5, we have $\left\lVert r\right\rVert_{1}\leq\epsilon N+6\epsilon N=7\epsilon N$. For the term I, we have
@@
By the standard concentration bounds (see A.6 and A.7), with high probability we have
@@
For $\text{I}_{2}$, by the argument above, Lemma B.8, and Corollary B.10 with appropriate choice of $\epsilon$, with high probability we have
@@
Again by Lemma B.8 and Corollary B.10, with high probability we have
@@
Bound on II. By Lemma B.12 (recall the definition of $\gamma_{2}$ (B.13) and the fact that $N>k^{\star}$ in this case) or B.14, with high probability we have
@@
Bound on III. By Lemma B.16, with high probability we have
@@
Substituting $\nu$ with $P^{\star}\mu$, the key is to require each term on the right-hand-side in (B.23) to be asymptotically smaller than the term $N^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}$. For convenience, we rewrite the definition of $\beta_{1}$ and $\beta_{2}$ here.

### v3

@@ 11  Set $\lambda=\left\lVert Y^{\top}D(\omega)Y-N\mathbf{I}_{k^{\star}}\right\rVert_{2}$
Set $\left\lVert\nu\right\rVert_{2}=0$ in Corollary B.14, we have
$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)(\widetilde{Y}_{i}-\nu)^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq C_{1}\left(\sqrt{Nk^{\star}}+\sqrt{N\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)\right).$ (C.18)
By normalization and standard concentration tools (see B.6), we have
$\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)^{\top}\nu\right\rVert_{2}=\left\lVert\nu\right\rVert_{2}\cdot\left\lVert\sum\limits_{i=1}^{N}(\widetilde{Y}_{i}-\nu)\right\rVert_{2}=C_{0}\cdot\sqrt{Nk^{\star}}\cdot\left\lVert\frac{1}{\sqrt{N}}\sum\limits_{i=1}^{N}\frac{\widetilde{Y}_{i}-\nu}{\sqrt{k^{\star}}}\right\rVert_{2}\leq c_{2}\sqrt{Nk^{\star}}$ (C.19)
with (high) constant probability. Now combine (C.18) and (C.19) to get
@@
Bound on II. Since upon the finish the algorithm, we have $\left\lVert\mathbf{1}-\omega\right\rVert_{1}\leq 2\epsilon N$, by Corollary B.16, we know
@@
Bound on III. Since $\left|S\cup C\right|\leq 2\epsilon N$, by Corollary B.16 and an argument similar to the one in II, we also have
@@
$\left\lVert\sum\limits_{i\in[N]\backslash(S\cup C)}\omega_{i}Y_{i}^{\top}Y_{i}-\left|[N]\backslash(S\cup C)\right|\nu\nu^{\top}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}\leq 4c\gamma_{2}.$ (C.20)
Now combine (C.17) and (C.20) and use the triangular inequality, we find that
@@
Again, by choosing the constant in the definition of $\gamma_{2}$ (C.13) sufficiently large, we can make $c$ above sufficiently small. Therefore, we can assume $c<\frac{1}{4}$, which leads to $\left\lVert\sum\limits_{i=1}^{N}\omega_{i}Y_{i}^{\top}Y_{i}-N\mathbf{I}_{k^{\star}}\right\rVert_{2}>\gamma_{2}$, a contradiction of the control condition of Algorithm 10. The proof is completed. ∎
@@
$\gamma_{3}=c\left(\sqrt{Nk^{\star}}+\sqrt{k^{\star}\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)+\epsilon N\ln\left(\frac{1}{\epsilon}\right)\right)$ (C.21)
@@
$\omega_{i}=\left(1-\frac{\tau_{i}}{\max\limits_{i}\tau_{i}}\right)\omega_{i},1\leq i\leq N.$ (C.22)
@@
1Set $\gamma_{3}$ as (C.21), and $\lambda=\left\lVert\sqrt{D(\omega)}\mathbf{Y}\mathbf{Y}^{\top}\sqrt{D(\omega)}-D(\omega)\right\rVert_{2}$. ($\omega$ is initialized as $\mathbf{1}$.)
@@ 4  Compute $\tau_{i}=\frac{v_{i}^{2}}{\omega_{i}}\mathbf{1}_{\left\{w_{i}>0\right\}}$
Proof of (i) is same as the one in the proof of Lemma C.12.
@@
Now assume $\omega^{(t)}\in\Lambda_{2}$ and we are still in the iteration, we will prove $\omega^{(t+1)}\in\Lambda_{2}$. By the definition, we only need to prove $5\sum\limits_{i\in C}v_{i}^{2}\geq\sum\limits_{i\in[N]\backslash C}v_{i}^{2}$. Suppose this does not hold, then $\left\lVert v_{C}\right\rVert_{2}^{2}<\frac{1}{6}$. By Corollary B.14 (set $\left\lVert\nu\right\rVert_{2}=0$), we have $\left\lVert\sqrt{D(\omega_{[N]\backslash C}^{(t)})}Y_{[N]\backslash C}Y_{[N]\backslash C}^{{}^{\top}}\sqrt{D(\omega_{[N]\backslash C}^{(t)})}-D(\omega_{[N]\backslash C}^{(t)})\right\rVert_{2}\leq c\gamma_{3}\leq c\lambda^{(t)}$ with high probability. However, since $v$ is the eigenvector corresponding to the maximal eigenvalue, we have
@@
Noting that $\left\lVert v_{C}\right\rVert_{2}^{2}+\left\lVert v_{[N]\backslash C}\right\rVert_{2}^{2}=1$, define $f_{c}(x)=c+(1-c)x+2\sqrt{x(1-x)},0\leq x<\frac{1}{6}$. Note that by selecting the constant in (C.21) sufficiently large, we can let $c\rightarrow 0$. It can be proved that in this case $0\leq f_{0}(x)<1$, which is a contradiction. (ii) is proved.
@@
After Algorithms 10 or 11, we have filtered the samples to ensure the second condition of $\omega$-regularity. Now we turn to the check of the final condition.
Actually, the weight $\omega$ output by Algorithms 10 or 11 already provides us with some useful insight.
Lemma C.15.
Set $\beta_{1}$, $\beta_{2}$ according to (C.11) and $\gamma_{i},i\in\left\{2,3\right\}$ according to (C.13) or (C.21) depending on whether $N\geq k^{\star}$ or not. For the weight $\omega$ output by the Algorithms 10 or 11 and any S[$\CS\subset[N]\backslash$ C with $\left|S\right|\leq\epsilon N$, we have
@@
By Lemma C.8, with high probability the original samples $\widetilde{\mathbf{Y}}$ are $(2\epsilon,\beta_{1},\beta_{2})$-regular. By spliting $N]=([N]\backslash C)\cup C$, we have
@@
Bound on $\text{I}_{1}$. Because $\left|S\right|\leq\epsilon N$, from $\omega$-regularity and Corollary C.10, we know $\text{I}_{1}\leq c_{1}\sqrt{N}\beta_{1}$.
Bound on $\text{I}_{2}$. Because $\omega$ is output by the Algorithms 10 or 11, we have $\left\lVert 1-\sqrt{\omega}\right\rVert_{1}\leq\left\lVert 1-\omega\right\rVert_{1}\leq\epsilon N$. By Corollary C.10 again, we have $\text{I}_{2}\leq c_{2}\beta_{2}$.
Bound on II. From Corollary C.10 and the fact $S\cap C=\varnothing$, we know $\text{II}\leq c_{3}\beta_{2}$.
Bound on III. From Lemma C.12, C.14, and the polarization equality, we have $\text{III}\leq c_{4}\epsilon N\gamma$.
@@
We are now prepared for the final algorithm, which is designed to ensure the third condition of $\omega$-regularity. Given samples $S$ and weight $\omega$ by the algorithms 10 or 11, we compute $\tau_{i}=\left|\left\langle\sqrt{\omega_{i}}Y_{i},\sum\limits_{j=1}^{N}\sqrt{\omega_{j}}Y_{j}\right\rangle-\omega_{i}k^{\star}\right|\mathbf{1}_{\left\{\omega_{i}>0\right\}}$. Then, we set $\omega_{i}$ to zero for the indices $i$ corresponding to the $\epsilon N$ maximal $\tau_{i}$.
@@
This lemma is very similar to Lemma C.15 but stronger in the sense that we allow any $S\subset[N]$ rather than S[$\CS\subset[N]\backslash$ C. This advantage is exactly granted by the sorting and pruning processes in Algorithm 12.
@@
The proof of this lemma is not hard but depends on the following fact. Because we have proved in Lemma C.15 that for any S[$\CS\subset[N]\backslash$ C and $\left|S\right|\leq\epsilon N$, we have $\left|\sum\limits_{i\in S}\sum\limits_{j=1}^{N}\sqrt{\omega_{i}\omega_{j}}\left\langle Y_{i},Y_{j}\right\rangle-k^{\star}\left\lVert\omega_{S}\right\rVert_{1}\right|\leq c(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma)$, we know that this bound also holds for any $S\subset[N]\backslash\left\{1,2,\dots,\epsilon N\right\}$. With the proning on $\omega$ to remove the first $\epsilon N$ coordinates, we can assume without loss of generality that $S\subset[N]\backslash\left\{1,2,\dots\epsilon N\right\}$. We have
@@
We have finished the proofs of Lemma C.11, C.12, C.14, and C.16, which means we have verified the correctness of $\omega$-regularity for our samples with high probability. The final step is to exploit the property of $\omega$-regularity to distinguish between $H_{0}$ and $H_{1}$.
Lemma C.17.
Let $\omega$ be the weights after running through the algorithm 9, 10 or 11, and 12, and $\beta_{1},\beta_{2}$ seleced as in Lemma C.8. Then, with high probability we have
$\left|\left\lVert\sum\limits_{i=1}^{N}\sqrt{\omega_{i}}Y_{i}\right\rVert_{2}^{2}-k^{\star}\left\lVert\omega\right\rVert_{1}-N^{2}\left\lVert\nu\right\rVert_{2}^{2}\right|\leq c\left(\sqrt{N}\beta_{1}+\beta_{2}+\epsilon N\gamma+\underbrace{\left(N^{\frac{3}{2}}\left\lVert\nu\right\rVert_{2}+\sqrt{Nk^{\star}}\right)\sqrt{\ln\left(\frac{1}{\alpha}\right)}+\ln\left(\frac{1}{\alpha}\right)}_{:=\text{R}}\right).$ (C.23)
@@
Bound on I. Define the vector $r\in\mathbb{R}^{N}$ where $r_{i}=1$ for $i\in C$ and $1-\sqrt{\omega_{i}}$ for i[$\Ci\in[N]\backslash$ C. By Algorithm 10 or 11, we have $\left\lVert r\right\rVert_{1}\leq\epsilon N+6\epsilon N=7\epsilon N$. For the term I, we have
@@
By the standard concentration bounds (see B.6 and B.7), with high probability we have
@@
For $\text{I}_{2}$, by the argument above, Lemma C.8, and Corollary C.10 with appropriate choice of $\epsilon$, with high probability we have
@@
Again by Lemma C.8 and Corollary C.10, with high probability we have
@@
Bound on II. By Lemma C.12 (recall the definition of $\gamma_{2}$ (C.13) and the fact that $N>k^{\star}$ in this case) or C.14, with high probability we have
@@
Bound on III. By Lemma C.16, with high probability we have
@@
Substituting $\nu$ with $P^{\star}\mu$, the key is to require each term on the right-hand-side in (C.23) to be asymptotically smaller than the term $N^{2}\left\lVert P^{\star}\mu\right\rVert_{2}^{2}$. For convenience, we rewrite the definition of $\beta_{1}$ and $\beta_{2}$ here.